How many moles of H2O will be produced from 2.9 grams of HCl?



1
Expert's answer
2022-02-16T11:17:03-0500

It actually depends on the reaction that takes place. But, in general, assuming that hydrogen atoms of HCl are the only ones to be converted to H2O, the reaction will look like

X + 2HCl --> Y + H2O


"n(H_2O)=2.9\ g(HCl)\times\frac{1\ mol(HCl)}{36.46\ g(HCl)}\times\frac{1\ mol(H_2O)}{2\ mol(HCl)}=0.040\ mol"


Again, this is just the possible scenario. There is not enough information to solve the problem precisely.